OpenStudy (anonymous):

7 sides A) two thousand eight hundred eighty degrees 12 sides B) nine hundred degrees 14 sides C) one thousand eight hundred degrees 18 sides D) two thousand one hundred sixty degrees match please very important

OpenStudy (campbell_st):

well the formula for angle sum of a polygon is \[Sum = ( n - 2) \times 180\] n = the number of sides.... as an example a 7 sided polygon n = 7 \[Sum = (7 - 2) \times 180\] or \[Sum = 5 \times 180...... so...... Sum = 900\] use the formula for the onther 2 polygons